OpenStudy (anonymous):

expand the logarithm log4 (3xyz)^2 explain how you got your answer step by step

OpenStudy (anonymous):

Use the logarithm properties. \[\log(a \times b) = loga + logb\] \[\log(\frac{ a}{ b }) = loga - logb \] \[loga ^{x} = xloga\]

OpenStudy (anonymous):

so can you do that with the equation please

OpenStudy (anonymous):

\[6logx + 2logy + 2logz\] I think..

OpenStudy (anonymous):

this is the answer i just need the steps to get it